• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name S. Kowalski

    Log entry time 15:58:16 on July27,2005

    Entry number 149682

    keyword=Shift Summary

    8:00 am Production running  
    8:55  Beam Off.  MCC Beam tests.  Pulsed beam.
          Stopped run
          Stopped Compton
    
    9:20  Beam back at 37 micro-A  Run 3867
          Started Compton Run
    9:30  Beam Off.  MCC forgot to do harp scan.  Stopped run and stopped Compton.
    
    10:55  Beam back.  5 micro-A.  No target.  Raster on. Run Spot++.
    
    10:58  He target back in.
    
    11:03  Beam back.  37 micro-A
           Started run 3868
    11:20  Beam off.  Run stopped.  Hall-C shutting off some power.
           Beamline magnet tripped.  Delay in restoring beam.
    
    11:52  Beam back, 37 micro-A.  Start run 3869.  Adjusting Compton chicane.  
    Turned Energy and position feedback back on.
    
    12:00  Lost left spectrometer control.  Lost IOC.  Rebooted.  Reset Q1 and Q2.  
    stopped run. Rest dipole currents.
    
    12:37  Start new run 3870. 
    
    14:36  Beam off.  More accelerator studies by MCC.  Target masked.  Helium 
    target out.